# On the Feasibility of Fingerprinting Collaborative Robot Traffic

### v2

@@ 2.1 Control Interfaces in Robotics
Script-based control. In this model, robots execute predefined scripts that automate tasks with minimal human intervention. The controller sends structured command messages and the robot executes them accordingly. For example, a domestic robot may be programmed to open curtains at a set time or sort recyclables from trash. These scripts, written in programming languages or proprietary software, define task logic that has evolved from simple sequences to complex behaviors influenced by perception, probabilistic decision-making, and machine learning. Commands are transmitted from the controller to the robot’s onboard computer, which executes actions and generates feedback. This feedback—such as vision, location, or object classification—is sent back to the controller and influences subsequent commands. For instance, in a pick-and-place task, the feedback helps determine grasp adjustments or movement corrections.
@@ 2.3 Encrypted Traffic Fingerprinting
While we now address earlier efforts aimed at fingerprinting robot traffic, we refer to Section 9, where we offer a broader perspective about other traffic fingerprinting contexts (e.g., website [37], video [33], and IoT [9] fingerprinting).
@@ 3.2 Experimental Setup
Robot actions dataset. For conducting our study, we gather a set of network traces generated when operating the robot to perform four different kinds of actions, including: a) picking up and placing an object, b) pouring water, c) turning on a switch, and d) pressing a key. We provide a depiction of the robotic arm while performing each of these actions in Figure 2. The actions share a certain degree of similarity; for instance, both pouring water and turning on a switch involve fluid motions and subtle changes in the robot’s articulation, which could be reflected in the traffic patterns as similar command message sequences, making it challenging to distinguish between them accurately. The duration of each trace also varies significantly across actions to capture cases where users might take different times to complete the same action. However, all our samples are bounded to a minimum and maximum completion time of 5 and 30 seconds.
We obtain our dataset by instructing the robot to execute the four actions, and collecting the network traffic exchanged between the robot and the controller. We collect a total of 200 samples: 50 for each action. The robot executes actions according to a set of predefined rules that an operator (ourselves) writes in scripts. However, data collection depends on the robot’s physical operation, which cannot be easily parallelized without additional robots (unavailable to us). The catalog of commands we execute includes different combinations of Cartesian motions (i.e., instructions to move the robot arm in a straight line along the X, Y, and Z axes), the opening/closing of a gripper claw, and adjustments of the gripper’s speed. This flexibility allows us to capture a multitude of real-world “trajectories”, where different operators may script the completion of an identical task with a different number of high-level commands or change their order. Each action sample follows a unique trajectory, composed of sub-tasks (e.g., movement, object manipulation) executed at varying points in time. Sub-tasks differ in speed, waypoints, object locations, and task-specific parameters (e.g., pouring angles), enhancing trajectory diversity. We leave the exploration of other approaches to generate command traffic sequences automatically (e.g., generative adversarial networks [42]) to future work.
Considered traffic analysis attacks. For showcasing the potential threats of traffic analysis attacks on robot control traffic, we leverage the open-source implementations of attacks used in the context of website fingerprinting, including CUMUL [29], k-FP [19], Tik-Tok [31], and Robust Fingerprinting (RF) [35]. These attacks rely either on manually-engineered traffic features, which make heavy use of summary statistics (CUMUL and k-FP), or on latent features extracted from network traces via deep learning architectures (Tik-Tok and RF).
@@ 3.3 Metrics
Defense evaluation metrics. When evaluating the defense mechanisms implemented to protect against traffic analysis, we primarily focus on two aspects: the reduction of an attack’s accuracy (effectiveness), and the impact of the defense on the robot’s performance (efficiency). We measure the efficiency of the defense through the bandwidth utilization and latency increase experienced by the robot’s activities. To compute bandwidth utilization overheads, we compare the amount of additional data exchanged due to the defense. To compute latency impacts, we compute the delays introduced in the communication between the robot and the controller, and assess the potential impacts of such delays on the correct operation of the robot and its ability to complete a designated task.
@@ 4 Characterization of Robot Actions
Individual commands have stable traffic signatures. Each row of Figure 3 shows two different network traces for a Cartesian command and a gripper movement command, respectively. We can see that different executions of the same command result in similar traffic patterns, but that these patterns also differ amongst each different kind of command. In more detail, Cartesian commands involve moving the robot arm in three-dimensional space and typically span a shorter period of time compared to gripper position commands. These Cartesian movements usually elicit a longer feedback packet from the robot arm, indicative of the arm’s positional adjustments in a 3D space. In contrast, gripper position command messages control the opening and closing of the robot’s gripper, and consist of packets just over 100 bytes in either direction. Each of these packets corresponds to an adjustment in the gripper’s position.
@@
Different actions result in different traffic patterns. Figure 4 shows example traffic patterns for the four classes of actions included in our dataset, showcasing both incoming and outgoing traffic from the point of view of the controller workstation. At a high level, we can see that each different action generates a disparate traffic pattern. For instance, the pick and place action generally involves at least two distinct gripper command messages, before and after the act of picking and placing an object. This is reflected in the traffic (Figure 2a) as two dense clusters of gripper position or speed command messages, indicating the gripper’s motion to open and close. Despite two actions being markedly dissimilar, the traces generated by the pour water action (Figure 2b) exhibit some similarities tied to specific commands that instruct the closing of the gripper to grab the bottle and opening it after pouring.
In turn, actions such as turning on a switch (Figure 2c) or pressing a key (Figure 2d) typically involve a tapping motion, which requires the robot to close the gripper beforehand. These actions do not necessitate additional gripper motions afterwards, although the gripper may open if required by subsequent tasks. The timing between Cartesian command messages also offers intricate details; tapping motions are usually swift, leading to shorter intervals between command messages. Other noteworthy detail includes the observation of interleaved occurrence of different commands. For instance, in the pick and place action, we can observe the presence of Cartesian command message between two gripper command messages, as seen in Figure 4c, around the 25 second mark. These temporal dependencies that are characteristic to each action may provide actionable information for building an effective classifier.
@@ 6.4 Evaluation of our Approach
Feature importance. The top 20 most important features for our classifier’s performance (depicted in Figure 9(c)) comprise a mix of signal processing and summary statistics elements. Specifically, from the signal processing category, the gripper speed average length emerges as the most significant feature, followed by the gripper speed command correlation coefficient median, ranking as the fourth most important, highlighting the role of gripper movement clusters in the classification process. On the other hand, from the summary statistics category, the total number of incoming packets and the 20th percentile of outgoing packet inter-arrival times stand out as the second and third most important features, respectively. This highlights that different classes exhibit variations not only in the total number of packets sent but also in the timing between these packets. Such distinctions enable our classifier to correctly distinguish between classes.
@@ 7.1 Considered Defenses
Padding packet sizes. This defense involves rounding up each packet size to the nearest hundredth (e.g., the nearest 200 bytes), and aims to obfuscate the actual size of the packets, making it more challenging for an adversary to discern patterns based on packet size. This method does not introduce additional latency, as it does not affect the transmission time of the packets. However, it does result in additional bandwidth overhead due to the transmission of larger-sized packets.
@@ 7.2 Evaluation of Defenses
Padding-only defense. Figure 10(a) illustrates the relationship between the accuracy of our classifier and the size of padded packets when the simpler padding-only defense is employed. The classifier retains a relatively high accuracy (above 80%) for packets padded until a maximum of 700 bytes, decreasing its accuracy to $\sim$40when packets are padded to 1000 bytes. For analyzing the security/efficiency trade-off in more detail, Figure 10(b) depicts the average bandwidth overhead percentage of the defense when considering our robot actions dataset. As observed, the overhead percentage increases linearly with the padding size, reaching approximately 700% overhead when the padding size is set to 800 bytes. At this padding size, the accuracy of our classifier is still relatively high ($\sim$70%). Thus, we conclude that this defense is largely ineffective for low padding sizes, achieving only a moderate protection for a large bandwidth overhead.
@@ 9 Related Work
Video fingerprinting. Video streaming protocols, such as Dynamic Adaptive Streaming over HTTP (DASH) or HTTP Live Streaming (HLS), which adjust their bit rates in response to varying network conditions, manifest traffic patterns that fluctuate based on the video content and its current quality or resolution. These fluctuations enable the development of video streaming fingerprinting techniques which can identify a given video being streamed, even if it is encrypted [33, 32]. Proposed defenses include the generation of synthetic video flows that shield the videos being watched by users [38], or noise addition [23].
@@ 10 Conclusion
In this paper, we investigated encrypted traffic analysis in robotics, emphasizing its implications for sensitive applications (e.g., domestic, healthcare). Leveraging machine learning classifiers fueled with signal-based processing features, we were able to discern robotic actions from encrypted traffic patterns, shedding light on significant privacy concerns. Our findings highlight potential vulnerabilities within robotic communication channels, calling for the development of enhanced security measures as the deployment of collaborative robots becomes widespread. Moreover, while we focused on collaborative robots, our method and findings likely also apply to industrial robot arms, such as those used in manufacturing. Their repetitive, pre-configured actions may make them even more vulnerable to traffic analysis attacks. While this study advances our understanding of the above issues, future work should focus on devising stronger and more efficient defense mechanisms in the realm of collaborative robotics security.

### v3

# On the Feasibility of Fingerprinting Collaborative Robot Network Traffic

@@ 2.1 Control Interfaces in Robotics
Script-based control. In this model, robots execute predefined scripts that automate tasks with minimal human intervention. The controller sends structured command messages, and the robot executes them accordingly. For example, a domestic robot may be programmed to open curtains at a set time or sort recyclables from trash. These scripts, written in programming languages or proprietary software, define task logic that has evolved from simple sequences to complex behaviors influenced by perception, probabilistic decision-making, and machine learning. Commands are transmitted from the controller to the robot’s onboard computer, which executes actions and generates feedback. This feedback—such as vision, location, or object classification—is sent back to the controller and influences subsequent commands. For instance, in a pick-and-place task, the feedback helps determine grasp adjustments or movement corrections.
@@ 2.3 Encrypted Traffic Fingerprinting
While we now address earlier efforts aimed at fingerprinting robot traffic, we refer to Section 9 for a broader perspective about other traffic fingerprinting contexts (e.g., website [37], video [33], and IoT [9] fingerprinting).
@@ 3.2 Experimental Setup
Robot actions dataset. We obtain our dataset by collecting the network traces exchanged between the robot and its controller when instructing the robot to execute four distinct actions: a) picking up and placing an object, b) pouring water, c) turning on a switch, and d) pressing a key. Figure 2 provides a depiction of the robotic arm while performing each of these actions. The actions share a certain degree of similarity; for instance, both pouring water and turning on a switch involve fluid motions and subtle changes in the robot’s articulation, which could be reflected in the traffic patterns as similar command message sequences, making it challenging to distinguish between them accurately. The duration of each trace also varies significantly across actions to capture cases where users might take different times to complete the same action. However, all our samples are bounded to a minimum and maximum completion time of 5 and 30 seconds.
We collect a total of 200 samples: 50 for each action. The robot executes actions according to a set of predefined rules that an operator (ourselves) writes in scripts. However, data collection depends on the robot’s physical operation, which cannot be easily parallelized without additional robots (unavailable to us). The catalog of commands we execute includes different combinations of Cartesian motions (i.e., instructions to move the robot arm in a straight line along the X, Y, and Z axes), the opening/closing of a gripper claw, and adjustments of the gripper’s speed. This flexibility allows us to capture a multitude of real-world trajectories, where different operators may script the completion of an identical task with a different number of high-level commands or change their order. Each action sample follows a unique trajectory, composed of sub-tasks (e.g., movement, object manipulation) executed at varying points in time. Sub-tasks differ in speed, waypoints, object locations, and task-specific parameters (e.g., pouring angles), enhancing trajectory diversity. We leave the exploration of other approaches to generate command traffic sequences automatically (e.g., generative adversarial networks [42]) to future work.
Considered traffic analysis attacks. For showcasing the potential threats of traffic analysis attacks on robot traffic, we leverage the open-source implementations of attacks used in the context of website fingerprinting (Section 5), including CUMUL [29], k-FP [19], Tik-Tok [31], and Robust Fingerprinting (RF) [35]. These attacks rely either on manually-engineered traffic features (CUMUL, k-FP), or latent features extracted via deep learning (Tik-Tok, RF).
Considered traffic analysis defenses. We consider two defenses inspired by constant-rate padding approaches [5] (Section 7). The first involves padding individual packets to enforce uniform packet lengths. The second transmits fixed-size packets at a constant rate, potentially including dummy packets, while ensuring compliance with the scheduling constraints required for robot communication.
@@ 3.3 Metrics
Defense evaluation metrics. We primarily focus on two aspects: the reduction of an attack’s accuracy (effectiveness), and the impact of the defense on the robot’s performance (efficiency). We measure the efficiency of the defense through the bandwidth utilization and latency increase experienced by the robot’s activities. To compute bandwidth utilization overheads, we compare the amount of additional data exchanged due to the defense. To compute latency impacts, we compute the delays introduced in the communication between the robot and the controller, and assess the potential impacts of such delays on the correct operation of the robot and its ability to complete a designated task.
@@ 4 Characterization of Robot Actions
Individual commands have stable traffic signatures. Figure 3 shows two different network traces for a Cartesian command and a gripper movement command, respectively. We can see that different executions of the same command result in similar traffic patterns, but that these patterns also differ amongst each different kind of command. In more detail, Cartesian commands involve moving the robot arm in three-dimensional space and typically span a shorter period of time compared to gripper position commands. These Cartesian movements usually elicit a longer feedback packet from the robot arm, indicative of the arm’s positional adjustments in a 3D space. In contrast, gripper position command messages control the opening and closing of the robot’s gripper, and consist of packets just over 100 bytes in either direction. Each of these packets corresponds to an adjustment in the gripper’s position.
@@
Different actions result in different traffic patterns. Figure 4 shows example traffic patterns for the four classes of actions included in our dataset, showcasing both incoming and outgoing traffic from the point of view of the controller workstation. At a high level, we can see that each different action generates a disparate traffic pattern. For instance, the pick and place action (Figure 2a) generally involves at least two distinct gripper command messages, before and after the act of picking and placing an object. This is reflected in the traffic as two dense clusters of gripper position or speed command messages, indicating the gripper’s motion to open and close. Despite the two actions being markedly dissimilar, the traces generated by the pour water action (Figure 2b) exhibit some similarities tied to specific commands that instruct the closing of the gripper to grab the bottle and opening it after pouring.
In turn, actions such as turning on a switch (Figure 2c) or pressing a key (Figure 2d) typically involve a tapping motion, which requires the robot to close the gripper beforehand. These actions do not necessitate additional gripper motions afterwards, although the gripper may open if required by subsequent tasks. The timing between Cartesian command messages also offers intricate details; tapping motions are usually swift, leading to shorter intervals between command messages. Other noteworthy detail includes the observation of interleaved occurrence of different commands. For instance, in the pick and place action, we can observe the presence of Cartesian command message between two gripper command messages, as seen in Figure 4c around the 25 second mark. These temporal dependencies that are characteristic to each action may provide actionable information for building an effective classifier.
@@ 6.4 Evaluation of our Approach
Feature importance. The top 20 most important features for our classifier’s performance (depicted in Figure 9(c)) comprise a mix of signal processing and summary statistics elements. Feature importance is determined based on the average gain, which measures the improvement in the loss function when a feature is used for splitting in the XGBoost model. From the signal processing category, the gripper speed average length emerges as the most significant feature, followed by the gripper speed command correlation coefficient median, ranking as the fourth most important, highlighting the role of gripper movement clusters in the classification. On the other hand, from the summary statistics category, the total number of incoming packets and the 20th percentile of outgoing packet inter-arrival times stand out as the second and third most important features, respectively. This highlights that different classes exhibit variations not only in the total number of packets sent but also in the timing between these packets. Such distinctions enable our classifier to correctly distinguish between classes.
@@ 7.1 Considered Defenses
Padding packet sizes. This defense obfuscates the size of packets, making it more challenging for an adversary to discern patterns based on packet size. This method does not introduce additional latency, as it does not affect the transmission time of the packets. However, it does result in additional bandwidth overhead due to the transmission of larger-sized packets. The defense mechanism rounds up each packet size to the nearest multiple of $x$ 100 bytes. We assume $x$ to be an integer ranging from 1 to 10 since the maximum packet size observed in our traces was under 1000 bytes. We also assume padding to occur only up to the maximum MTU size of 1500 bytes. For example, a packet of size 360 bytes would be padded to 400 bytes if $x=2$ (nearest multiple of 200) and to 500 bytes if $x=5$, while a packet of size 960 bytes would be padded to 1500 bytes if $x=8$.
@@ 7.2 Evaluation of Defenses
Padding-only defense. Figure 10(a) illustrates the relationship between the accuracy of our classifier and the size of padded packets when the simpler padding-only defense is employed. The classifier retains a relatively high accuracy (above 80%) for packets padded to the nearest multiple of 700 bytes ($x=7$), decreasing its accuracy to $\sim$40when packets are padded to the nearest multiple of 1000 bytes ($x=10$). For analyzing the security/efficiency trade-off in more detail, Figure 10(b) depicts the average bandwidth overhead percentage of the defense when considering our robot actions dataset. As observed, the overhead percentage increases linearly with the padding size, reaching approximately 700% overhead when the padding size is set to the nearest multiple of 800 bytes ($x=8$). At this padding size, the accuracy of our classifier is still relatively high ($\sim$70%). Thus, we conclude that this defense is largely ineffective for low padding sizes, achieving only a moderate protection for a large bandwidth overhead.
@@ 9 Related Work
Video fingerprinting. Video streaming protocols that adjust their bit rates in response to network conditions manifest traffic patterns that fluctuate based on the video content and its resolution, enabling the development of fingerprinting techniques which can identify a given video being streamed, even if it is encrypted [33, 32]. Proposed defenses include the generation of synthetic video flows that shield the videos being watched by users [38], or noise addition [23].
@@ 10 Conclusion
In this paper, we investigated encrypted traffic analysis in robotics, emphasizing its implications for sensitive applications (e.g., domestic, healthcare). Leveraging machine learning classifiers fueled with signal-based processing features, we were able to discern robotic actions from encrypted traffic patterns, shedding light on significant privacy concerns. Our findings highlight potential vulnerabilities within robotic communication channels, calling for the development of enhanced security measures as the deployment of collaborative robots becomes widespread. Moreover, while we focused on collaborative robots, our method and findings likely also apply to industrial robot arms, such as those used in manufacturing. Their repetitive, pre-configured actions may make them even more vulnerable to traffic analysis attacks, casting the need for more advanced defense mechanisms.